One company can clear a parcel of land in 25 hours. Another company requires 30 hours to do the job. If both are hired, how long with it take to clear the parcel working together?
Answer by stanbon(75887)

One company can clear a parcel of land in 25 hours. Another company requires 30 hours to do the job. If both are hired, how long with it take to clear the parcel working together?
1st company DATA:
time = 25 hr/job ; rate = 1/25 job/hr
-----
2nd company DATA:
time = 30 hr/job ; rate = 1/30 job/hr
-----------
together DATA:
rate = (1/25)+(1/30) = (30+25)/(25*30) = 55/(25*30) = 11/150 job/hr
------
Ans: time to do the job together = 150/11 = 13.64 hrs/job = 13 hr + 38 min
----------
Cheers,
Stan H.
-----------
-------